COLOR-CENTER LASER SPECTRUM OF THIOPHENE OBSERVED WITH THE PULSED SUPERSONIC JET EXPANSION TECHNIQUE

A. OKUMURA, \underline{M. ISHIGURO}, K. TANAKA, and T. TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Higashiku, Fukuoka 812-8581, Japan.

The high resolution infrared spectrum of thiophene in the CH-stretch region (3070-3140 cm^{−1}) has been observed with a color-center laser spectrometer combined with a pulsed supersonic jet expandion technique. Of 4 fundamental bands (a_{1};ν_{1}, ν_{2}, b_{1}; ν_{12}, ν_{13}) expected in this region, the ν_{2} band was analyzed to the confirmation of previous vibrational assignment. ^{a} ^{b} The rotational constants and band origin were derived from the observed spectra while the ground state constants were fixed at the microwave values. ^{c} The ν_{2} band was found to be overlapped by an a_{1} vibrational band, which is denoted as ν_{x}. The ν_{x} band presumably borrows the intensity from the ν_{2} band by Fermi resonance. The vibrational assignment of ν_{x} is not get clear, but the magnitude of the interaction between the ν_{2} and ν_{x} states is calculated to be 0.8 cm^{−1} from the analysis of the intensity ratio of these bands.

aT. D. Klots, R. D. Chirico, and W. B. Steele, Spectrochim. Acta. \underline{50A}, 765 (1994)

bA. A. El-Azhary, and R. H. Hilal, Spectrochim. Acta. \underline{53A}, 1365 (1997)

cB. Pankoke, K. M. T. Yamada, and G. Winnewisser, Z. Naturforsch. \underline{49a}, 1193 (1994)